If the rate of appearance of oxygen over change in time is 6.0*10^-5ms^-1 at a particular instance.what is the rate of dissappearance of Oxygen3 at the same time.
1
Expert's answer
2019-06-11T11:25:29-0400

Balanced chemical equation is:


"2O_3(g)\rightarrow 3O_2(g)"

"Rate = -\frac{1}{2}\times \frac{\Delta [O_3]}{\Delta t} = \frac{1}{3}\times \frac{\Delta [O_2]}{\Delta t}"

Express yje rate of dissaperatnce of O3 from this equation:


"-\frac{\Delta [O_3]}{\Delta t} = \frac{2}{3}\times \frac{\Delta [O_2]}{\Delta t} = \frac{2}{3}\times (6.0\times 10^{-5}\frac{M}{s})= 4.0\times 10^{-5} \frac{M}{s}"